OpenStudy (anonymous):

ALGEBRA 2 Solve for x: 2 over 5 plus 3 over 5 x equals quantity x plus 5 over 10 Solve for x: 7 over 8 minus 1 over x equals 3 over 4.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{2}{5} + \frac{3}{5}x=\frac{x+5}{10} }\) like this ?

OpenStudy (anonymous):

the first one the first fraction is right but the second one is 3 over and 5x

OpenStudy (anonymous):

that made no sense, lol i mean 5x under the fraction

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{2}{5} + \frac{3}{5x}=\frac{x+5}{10} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{2x}{5x} + \frac{3}{5x}=\frac{x+5}{10} }\) \(\LARGE\color{blue}{ \frac{2x+3}{5x}=\frac{x+5}{10} }\)

OpenStudy (solomonzelman):

then cross multiply

OpenStudy (anonymous):

20x+30=10x^2+10x^2 ?